OpenStudy (zairahernandez):

Shona is buying a rug for her room. Store A has the rug for $45 with a 10% discount. Store B has the same rug for $46 and is offering a $10 rebate. The sales tax is 6% on either purchase. If Shona only has $40 to spend, which store will she purchase the rug from, and how much will she have left over? Shona can afford the rug in store A. She will have $2.93 left over. Shona can afford the rug in store B. She will have $1.84 left over. Shona can afford the rug in store B. She will have $2.93 left over. Shona can afford the rug in store A. She will have $1.84 left over.

OpenStudy (zairahernandez):

@undeadknight26

undeadknight26 (undeadknight26):

Store A= 45 x .10 = 4.5 45 - 4.5 = 40.5 40.5 x .06 = 2.43 40.5 + 2.43 = 42.93 Store B = 46 - 10 = 36 36 x .06 = 2.16 36 + 2.16 = 38.16

OpenStudy (zairahernandez):

Yes But Is Is A or b or c or d?

undeadknight26 (undeadknight26):

Hmm...what's your "guess"?

OpenStudy (zairahernandez):

I Have no Idea I only got 10 minutes eft on that answer xc

OpenStudy (zairahernandez):

A.Shona can afford the rug in store A. She will have $2.93 left over. B.Shona can afford the rug in store B. She will have $1.84 left over. C.Shona can afford the rug in store B. She will have $2.93 left over. D.Shona can afford the rug in store A. She will have $1.84 left over.

undeadknight26 (undeadknight26):

LOL right well ill tell ya this...You can eliminate A and D because it def can't be store A...Since 40 - 42 is a negative number...So it has to be either B or C...And seeing this its 40 - 38.16

undeadknight26 (undeadknight26):

So the answer is B.

OpenStudy (anonymous):

i was right

OpenStudy (anonymous):

just was not sure

undeadknight26 (undeadknight26):

lol good job mate.

undeadknight26 (undeadknight26):

If ya need anything else....Well ask me tomorrow because my big brother is being mean ._.

OpenStudy (zairahernandez):

Thank you guys

OpenStudy (anonymous):

no thx u